How would you solve this equation?
11m^2-14m=0

Answer by stanbon(75887)

How would you solve this equation?
11m^2-14m=0
How would you solve this equation?
11m^2-14m=0
--------------
m(11m-14) = 0
m= 0 or m = 14/11
-----------------------
Cheers,
Stan H.